OpenStudy (anonymous):

Evaluate 30 - a2 ÷ c + (11 - b) if a = 6, b = 7, c = 4

OpenStudy (callisto):

\[ 30 - a^2 ÷ c + (11 - b)\]\[=30 - 6^2 ÷ 4 + (11-7)\]\[=30 - 9 + 4=?\]

OpenStudy (anonymous):

thx

OpenStudy (callisto):

welcome

OpenStudy (anonymous):

but whats the answer